OpenStudy (anonymous):

-2/3 ^3 -2^2 4^3/3 evaluate

OpenStudy (anonymous):

Are there any parentheses in your problems?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

On the first one.. is the ^3 for the entire fraction?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The exponent on goes with the fraction. - 2 * 2 * 2 8 -------- = - --- 3 * 3 * 3 27

OpenStudy (anonymous):

-2^2 would be - 2 * 2 = -4

OpenStudy (anonymous):

4^3 = 4 * 4 * 4 -------- 3 64 -- are you sure the problem is correct? 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then it is 64/3 or 21 1/3